OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (ddcamp):

First, find the slope: \[m = \frac{ y_2-y_1 }{ x_2-x_1 }\] Then, use that in the point slope equation: \[y-y_1=m(x-x_1)\] Then, solve for standard form.

OpenStudy (anonymous):

Thank you so so much @DDcamp